OpenStudy (jaylelile):

help me!!! find solutions to graph 2x-y=-2

OpenStudy (electronia):

-1

OpenStudy (electronia):

I think.

OpenStudy (electronia):

POSITIVE 1 for the first part actually.

OpenStudy (jaylelile):

x=-1 you say? wait what?? I'm trying to find ordered pairs so I can graph the line.

OpenStudy (electronia):

Oh, I thought you where solving the equation, did not read. Sorry bout that!

OpenStudy (jaylelile):

no problem lol

OpenStudy (trojanpoem):

This equation has infinite number of solutions. However, from your comment I think you're trying to graph it , right ?

OpenStudy (jaylelile):

yes

OpenStudy (jaylelile):

I just need 3 solutions and im unsure how to find them

OpenStudy (jaylelile):

@TrojanPoem

OpenStudy (thadds2003):

do you still need help?

OpenStudy (thadds2003):

ok then.

OpenStudy (okdutchman7):

It can be simplified to 2x + 2 Giving these values for (x,y) (-10,-18) (-9,-16) (-8,-14) (-7,-12) (-6,-10) (-5,-8) (-4, -6) (-3, -4) (-2, -2) (-1, 0) (0, 2) (1, 4) (2, 6) (3, 8) (4, 10) (5, 12) (6, 14) (7, 16) (8, 18) (9, 20) (10, 22)

OpenStudy (jaylelile):

but I still don't know how you got that....

OpenStudy (trojanpoem):

To Graph the function: - Isolate the variable y. (y = 2x+2) - Pick any three points of your choice. (Personally I pick -1, 0, 1 and if I didn't get the absolute shape I would pick -2 , 2 ... until I fully understand it) when x = -1 y = -2 + 2 = 0 when x = 0 y = 2 when x = 1 y = 4 

OpenStudy (trojanpoem):

y = 2x + 2 This is a first degree function. so It will be a straight line y = mx + b b : y - intercept. m : slope of the line.

OpenStudy (trojanpoem):